# Learning Joint ID-Textual Representation for ID-Preserving Image Synthesis

### v1

@@ Abstract
Recent advancements in text-to-image (T2I) diffusion models have significantly enhanced the quality and versatility of image generation. A persistent challenge in personalized image synthesis is maintaining the subject’s identity (ID) from reference images while allowing flexible modification of other visual attributes. In this paper, we propose a novel framework for ID-preserving image generation. Instead of injecting identity features via an adapter module into a pre-trained T2I model, we achieve both identity preservation and text alignment through a multi-modal encoding strategy. Specifically, we treat identity and textual semantics as a unified prior condition to guide the generation process. To this end, we introduce FaceCLIP, a multi-modal encoder that learns a joint embedding space for both human identity and textual semantics. Given a reference face image and a user-provided text prompt, FaceCLIP produces a joint representation that encapsulates both identity and text information. This unified representation is then used to condition a base generative model, enabling it to synthesize portrait images that are both identity-consistent with the reference image and semantically aligned with the text prompt. We further develop a multi-modal alignment algorithm to train FaceCLIP. The encoder is optimized using a multi-modal alignment loss that explicitly aligns its joint representations with the original face embedding space, text embedding space, and image embedding space. Building on this, we construct a new ID-preserving image synthesis pipeline, FaceCLIP-SDXL, by integrating FaceCLIP with Stable Diffusion XL (SDXL). Compared with the existing ID-preserving image synthesis methods, our methods allows photorealistic portrait image generation with superior ID preservation capability and text adherence. The extensive experiments show that FaceCLIP-SDXL outputs the existing methods quantitatively and qualitatively.
@@ 3.1 Problem Formulation
We formulate the data distribution of portrait images as a conditional distribution based on the subject’s identity and textual description, denoted as $p(x\mid e)$. Here, $c$ represents a unified representation of both identity and text conditions, defined as:
@@
$L_{\text{DM}}=\mathbb{E}_{x_{0},\epsilon,e\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\left\|\epsilon-\epsilon_{\phi}(x_{t},t,e)\right\|^{2}_{2}\right],$ (2)
@@ 3.2 Learning Joint Identity-text Encoding
where $e_{c\to t}\in\mathbb{R}^{b\times d_{t}}$ and $e_{c\to r}\in\mathbb{R}^{b\times d_{r}}$ are linear projections of the joint embedding $e$, used to match the dimensionality of image and ID embedding spaces, respectively. The contrastive loss function $\mathcal{L}_{c}(\cdot,\cdot)$ is implemented as shown in the pseudocode below:
1# x1 [b, d]: projected embedding batch
2# x2 [b, d]: projected embedding batch
3import numpy as np
4
5def contrastive_loss(x1, x2):
6 logits = np.dot(x1, x2.T)
7 loss1 = cross_entropy_loss(logits, axis=0)
8 loss2 = cross_entropy_loss(logits, axis=1)
9 loss = (loss1 + loss2) / 2
10 return loss
@@ 3.3 Pre-training Workflow
The high-level architecture of our FaceCLIP encoder $\mathcal{H}$ and training workflow is illustrated in Figure 1. The FaceCLIP encoder processes a batch of aligned face images $c_{r}\in\mathbb{R}^{b\times h_{r}\times w_{r}\times d_{r}}$ and text prompts $c_{t}$ as input, outputting fused embeddings $e$. Here, $h_{<*>}$ and $w_{<*>}$ denotes the spatial size of an image. Within the encoder, the face image is processed by a face encoder into a face class embedding $e_{r_{\text{cls}}}\in\mathbb{R}^{b\times d_{r_{\text{cls}}}}$ and face patch features $e_{r_{\text{pat}}}\in\mathbb{R}^{b\times L_{r}\times d}$. Meanwhile, the text prompt is converted into text class embeddings $e_{t_{\text{cls}}}\in\mathbb{R}^{d_{t_{\text{cls}}}}$ and text patch embeddings $e_{t_{\text{pat}}}\in\mathbb{R}^{b\times L\times d}$. The fused representation $e$ is obtained by feeding $e_{r_{\text{pat}}}$ and $e_{t_{\text{pat}}}$ into the Fusion Module. As defined in the previous section, $e$ serves as a unified representation of both identity and textual information. To enable the FaceCLIP encoder to learn a joint identity-text embedding space, we extend the vision-language pretraining method proposed in [42] with an additional identity alignment loss. As depicted in Figure 1, $e$ is further projected into projected face embeddings $e_{c\to r}$ and projected text embeddings $e_{c\to t}$. Similar to CLIP, we align $e_{c\to t}$ with image embeddings $e_{I}$ extracted from target images $x_{0}$ by a CLIP vision encoder using a contrastive loss [43]. However, applying contrastive loss solely on $e_{c\to t}$ and $e_{I}$ may lead to trivial text-image alignment. To address this, we introduce an additional contrastive loss to explicitly align $e_{c\to r}$ with $e_{r_{\text{cls}}}$, ensuring that $e$ preserves identity information. During training, the face encoder, CLIP text encoder, CLIP vision encoder, and the image projection layer are freeze.
@@ 3.4 Fusion Module
The Fusion Module (FM) is the key component responsible for integrating identity features and text features. It consists of multiple cascaded Feature Fusion (FF) blocks, with the architecture of a single FF block depicted in Figure 2. Each FF block takes text patch embeddings $e_{t_{\text{pat}}}$ and face patch embeddings $e_{r_{\text{pat}}}$ as input and outputs fused text embeddings $\hat{e}_{t_{\text{pat}}}$ and fused face embeddings $\hat{e}_{r_{\text{pat}}}$. Within each FF block, $e_{t_{\text{pat}}}$ and $e_{r_{\text{pat}}}$ first pass through a Dual Cross-Attention (DCA) module, which integrates text and identity information using a cross-attention mechanism. The mixed embeddings are then processed by two separate self-attention (SA) layers, yielding the final fused text embeddings $\hat{e}_{t_{\text{pat}}}$ and fused face embeddings $\hat{e}_{r_{\text{pat}}}$. Finally, the fused text embeddings from the last FF block serve as the joint identity-text representation $e$.
Figure 1: FaceCLIP architecture and pre-training workflow. The modules labeled in red are freezed and the modules labeled in blue are unfreezed during pre-training.
@@ 3.6 Training Diffusion Model with Joint Identity-text Priors
Figure 2: Fusion Module: (a) Architecture of Feature Fusion Block in Fusion Module; (b) Detailed computation graph of Dual Cross-Attention.
@@ 4 Experiments
We conduct a series of experiments to evaluate the proposed method from multiple perspectives: (1) to validate that the FaceCLIP encoder learns distinctive representations for both identity and textual semantics; (2) to assess the effectiveness of the FaceCLIP representation and pre-training strategy in enabling ID-preserving image generation; (3) to evaluate the overall performance of FaceCLIP-SDXL by comparing it with existing ID preservation; (4) to evaluate the performance of face synthesis by comparing FaceCLIP-SD1.5 with existing face synthesis models. Specifically, the text alignment performance is assessed via zero-shot classification accuracy on ImageNet-1k [8]. The quality of identity alignment is examined by visualizing the fused embedding $e$ using t-SNE [44]. Additionally, we conduct an ablation study to investigate the impact of the FaceCLIP representation on image generation quality. Finally, we compare our method against existing ID preservation and face synthesis approaches through both quantitative evaluations and qualitative assessments.
@@ 4.1 Setting
Implementation Details. We implement a FaceCLIP-L-14, a FaceCLIP-bigG-14, FaceCLIP-SD1.5 and FaceCLIP-SDXL for all experiments using PyTorch. All models are trained in a distributed manner using PyTorch’s built-in Distributed Data Parallel (DDP) library with numerical precision set to $bf16$. FaceCLIP-L-14 and FaceCLIP-bigG-14 are trained on 64 NVIDIA H100 GPUs with batch sizes of 24K and 16K, respectively. FaceCLIP-SD.15 and FaceCLIP-SDXL is trained on 128 NVIDIA H100 GPUs with a total batch size of 4096 and 512. To mitigate weight stagnation caused by $bf16$, we adopt the AnyPrecisionAdamW optimizer [45], setting $\beta$ to $(0.9,0.999)$ and weight decay to 0.01. The learning rate for all models is configured as $52e-$ 5.
Datasets. We construct a large-scale Single-Per-Single-View (SPSV) dataset comprising tuples of images, text descriptions, and reference images to train FaceCLIP encoder, FaceCLIP-SD1.5, and FaceCLIP-SDXL. The reference images are aligned face crops with a spatial size of $448\times 448$. Our dataset includes nine public datasets [46, 47, 48, 49, 50, 51, 52, 53, 54] alongside internal datasets. After rigorous preprocessing and filtering, we obtain a total of 43 million data samples. To preserve text alignment capability, we use an internal large-scale image-text dataset as guided data. We evaluate image synthesis performance on the Internal-v1 validation set and Unsplash-50 [11]. Internal-v1 includes 15 identities, each associated with 200 text prompts describing variations in location, weather, lighting conditions, and subject behavior. Unsplash-50 consists of 50 high-resolution portrait images with corresponding descriptions.
@@ 4.2 Text Alignment Verification
Model Top1 / Top5
OpenCLIP-L-14 75.2 / 94.3
OpenCLIP-bigG-14 80.1 / 96.0
FaceCLIP-L-14 45.8 / 64.2
FaceCLIP-bigG-14 47.2 / 69.9
FaceCLIP-L-14 75.3 / 94.9
FaceCLIP-bigG-14 76.9 / 95.1
@@ 4.4 Identity-Preserved T2I
4.4.1 Face Synthesis
Quantitative Assessment. To ensure a fair comparison with Arc2Face, which is developed based on Stable Diffusion 1.5 (SD1.5), we compare our downgraded variant, FaceCLIP-SD1.5, against Arc2Face in terms of face similarity and image fidelity, using both quantitative and qualitative evaluations. As shown in Table 2, FaceCLIP-SD1.5 outperforms Arc2Face, achieving higher face similarity scores of 0.930 on Internal-v1 and 0.919 on Unsplash-50. Moreover, FaceCLIP-SD1.5 demonstrates superior image quality, reflected by lower FID scores of 5.823 and 6.012, respectively. Qualitative comparisons are illustrated in Figure 4, where FaceCLIP-SD1.5 generates images with enhanced realism, improved fidelity, and more natural tonal rendering, while maintaining strong identity consistency.
Model Internal-v1 Unsplash-50
Face Sim. $\uparrow$ FID $\downarrow$ Face Sim. $\uparrow$ FID $\downarrow$
Arc2Face 0.914 7.313 0.897 7.624
FaceCLIP-SD1.5 0.930 5.823 0.919 6.012
Table 2: Comparing FaceCLIP-SD1.5 with Arc2Face in terms of face synthesis.
User Study. We further conducted a user study to compare FaceCLIP-SD1.5 with Arc2Face, focusing on face similarity and image fidelity. The selection rate in favor of FaceCLIP-SD1.5 reached 74.2%, compared to 25.8% for Arc2Face, indicating a clear human preference for our method in terms of perceived quality and identity resemblance.
Figure 4: Face images by Arc2Face and FaceCLIP-SD1.5.
@@ 4.4.2 ID Preservation Comparison
4.5 Ablation Study.
To evaluate the impact of our pre-training methodology, we conduct an ablation study by comparing two training variants of FaceCLIP-SDXL. We first pre-train the FaceCLIP-L-14 and FaceCLIP-bigG-14 encoders, and then apply them to train two FaceCLIP-SDXL variants. The first variant, FaceCLIP-SDXL, is trained by fine-tuning both the UNet and FaceCLIP’s fusion module. The second variant, FaceCLIP-SDXL, is trained by fine-tuning only the UNet, keeping the fusion module fixed. As shown in Table 4, fine-tuning the fusion module improves text alignment but slightly reduces face similarity. Image quality, measured by FID, remains largely unchanged across the two configurations. These results indicate that, under the encoder-plus-base-model architecture, training solely with diffusion loss tends to bias the model toward generic T2I synthesis objectives. This highlights the critical role of our pre-training strategy in enabling effective identity preservation.
Model Internal-v1 Unsplash-50
FaceCLIP-SDXL 0.082 0.359 87.1 0.093 0.368 88.1
FaceCLIP-SDXL 0.869 0.332 86.9 0.866 0.329 87.8
Table 4: Quantitative results of FaceCLIP-SDXL with different training strategies: $\spadesuit$ freezing entire FaceCLIP encoder after pre-training; $\clubsuit$ fine-tuning Fusion Block after pre-training.

### v3

@@ Abstract
Recent progress in text-to-image (T2I) diffusion models has greatly improved image quality and flexibility. However, a major challenge in personalized generation remains: preserving the subject’s identity (ID) while allowing diverse visual changes. We address this with a new framework for ID-preserving image generation. Instead of relying on adapter modules to inject identity features into pre-trained models, we propose a unified multi-modal encoding strategy that jointly captures identity and text information. Our method, called FaceCLIP, learns a shared embedding space for facial identity and textual semantics. Given a reference face image and a text prompt, FaceCLIP produces a joint representation that guides the generative model to synthesize images consistent with both the subject’s identity and the prompt. To train FaceCLIP, we introduce a multi-modal alignment loss that aligns features across face, text, and image domains. We then integrate FaceCLIP with Stable Diffusion XL, forming a complete synthesis pipeline named FaceCLIP-SDXL. Compared to existing ID-preserving approaches, our method produces more photorealistic portraits with better identity retention and text alignment. Extensive experiments demonstrate that FaceCLIP-SDXL outperforms prior methods in both qualitative and quantitative evaluations.
@@ 3.1 Problem Formulation
We formulate the data distribution of portrait images as a conditional distribution based on the subject’s identity and textual description, denoted as $p(x\mid e)$. Here, $e$ represents a unified representation of both identity and text conditions, defined as:
@@
$L_{\text{DM}}=\mathbb{E}_{x_{0},e,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t}\left[\left\|\epsilon-\epsilon_{\phi}(x_{t},t,e)\right\|^{2}_{2}\right],$ (2)
@@ 3.2 Learning Joint Identity-text Encoding
where $e_{c\to t}\in\mathbb{R}^{b\times d_{t}}$ and $e_{c\to r}\in\mathbb{R}^{b\times d_{r}}$ are linear projections of the joint embedding $e$, used to match the dimensionality of image and ID embedding spaces, respectively. $\mathcal{L}_{c}(\cdot,\cdot)$ is the contrastive loss function proposed in [42].
@@ 3.3 Pre-training Workflow
The high-level architecture of our FaceCLIP encoder $\mathcal{H}$ and training workflow is illustrated in Figure 1 (a). The FaceCLIP encoder processes a batch of aligned face images $c_{r}\in\mathbb{R}^{b\times h_{r}\times w_{r}\times d_{r}}$ and text prompts $c_{t}$ as input, outputting fused embeddings $e$. Here, $h_{<*>}$ and $w_{<*>}$ denotes the spatial size of an image. Within the encoder, the face image is processed by a face encoder into a face class embedding $e_{r_{\text{cls}}}\in\mathbb{R}^{b\times d_{r_{\text{cls}}}}$ and face patch features $e_{r_{\text{pat}}}\in\mathbb{R}^{b\times L_{r}\times d}$. Meanwhile, the text prompt is converted into text class embeddings $e_{t_{\text{cls}}}\in\mathbb{R}^{d_{t_{\text{cls}}}}$ and text patch embeddings $e_{t_{\text{pat}}}\in\mathbb{R}^{b\times L\times d}$. The fused representation $e$ is obtained by feeding $e_{r_{\text{pat}}}$ and $e_{t_{\text{pat}}}$ into the Fusion Module. As defined in the previous section, $e$ serves as a unified representation of both identity and textual information. To enable the FaceCLIP encoder to learn a joint identity-text embedding space, we extend the vision-language pretraining method proposed in [42] with an additional identity alignment loss. As depicted in Figure 1 (a), $e$ is further projected into projected face embeddings $e_{c\to r}$ and projected text embeddings $e_{c\to t}$. Similar to CLIP, we align $e_{c\to t}$ with image embeddings $e_{I}$ extracted from target images $x_{0}$ by a CLIP vision encoder using a contrastive loss [43]. However, applying contrastive loss solely on $e_{c\to t}$ and $e_{I}$ may lead to trivial text-image alignment. To address this, we introduce an additional contrastive loss to explicitly align $e_{c\to r}$ with $e_{r_{\text{cls}}}$, ensuring that $e$ preserves identity information. During training, the face encoder, CLIP text encoder, CLIP vision encoder, and the image projection layer are freeze.
@@ 3.4 Fusion Module
The Fusion Module (FM) is the key component responsible for integrating identity features and text features. It consists of multiple cascaded Feature Fusion (FF) blocks, with the architecture of a single FF block depicted in Figure 1 (b) and (c). Each FF block takes text patch embeddings $e_{t_{\text{pat}}}$ and face patch embeddings $e_{r_{\text{pat}}}$ as input and outputs fused text embeddings $\hat{e}_{t_{\text{pat}}}$ and fused face embeddings $\hat{e}_{r_{\text{pat}}}$. Within each FF block, $e_{t_{\text{pat}}}$ and $e_{r_{\text{pat}}}$ first pass through a Dual Cross-Attention (DCA) module, which integrates text and identity information using a cross-attention mechanism. The mixed embeddings are then processed by two separate self-attention (SA) layers, yielding the final fused text embeddings $\hat{e}_{t_{\text{pat}}}$ and fused face embeddings $\hat{e}_{r_{\text{pat}}}$. Finally, the fused text embeddings from the last FF block serve as the joint identity-text representation $e$.
Figure 1: (a) FaceCLIP architecture and pre-training workflow.The modules labeled in red are freezed and the modules labeled in blue are unfreezed during pre-training; (b) Architecture of Feature Fusion Block in Fusion Module; (c) Detailed computation graph of Dual Cross-Attention.
@@ 4 Experiments
We conduct a series of experiments to evaluate the proposed method from multiple perspectives: (1) to validate that the FaceCLIP encoder learns distinctive representations for both identity and textual semantics; (2) to assess the effectiveness of the FaceCLIP representation and pre-training strategy in enabling ID-preserving image generation; (3) to evaluate the overall performance of FaceCLIP-SDXL by comparing it with existing ID preservation. Specifically, the text alignment performance is assessed via zero-shot classification accuracy on ImageNet-1k [8]. The quality of identity alignment is examined by visualizing the fused embedding $e$ using t-SNE [44]. Additionally, we conduct an ablation study to investigate the impact of the FaceCLIP representation on image generation quality. Finally, we compare our method against existing ID preservation approaches through both quantitative evaluations and qualitative assessments.
@@ 4.1 Setting
Implementation Details. We implement a FaceCLIP-L-14, a FaceCLIP-bigG-14, and FaceCLIP-SDXL for all experiments using PyTorch. All models are trained in a distributed manner using PyTorch’s built-in Distributed Data Parallel (DDP) library with numerical precision set to $bf16$. FaceCLIP-L-14 and FaceCLIP-bigG-14 are trained on 64 NVIDIA H100 GPUs with batch sizes of 24K and 16K, respectively. FaceCLIP-SDXL is trained on 128 NVIDIA H100 GPUs with a total batch size of 4096 and 512. To mitigate weight stagnation caused by $bf16$, we adopt the AnyPrecisionAdamW optimizer [45], setting $\beta$ to $(0.9,0.999)$ and weight decay to 0.01. The learning rate for all models is configured as $52e-$ 5.
Datasets. We construct a large-scale Single-Per-Single-View (SPSV) dataset comprising tuples of images, text descriptions, and reference images to train FaceCLIP encoder and FaceCLIP-SDXL. The reference images are aligned face crops with a spatial size of $448\times 448$. Our dataset includes nine public datasets [46, 47, 48, 49, 50, 51, 52, 53, 54] alongside internal datasets. After rigorous preprocessing and filtering, we obtain a total of 43 million data samples with portrait images and captions annotated by InternVL [55]. To preserve text alignment capability, we use an internal large-scale image-text dataset as guided data. We evaluate image synthesis performance on the Internal-v1 validation set and Unsplash-50 [11]. Internal-v1 includes 15 identities, each associated with 200 text prompts describing variations in location, weather, lighting conditions, and subject behavior. Unsplash-50 consists of 50 high-resolution portrait images with corresponding descriptions.
@@ 4.2 Text Alignment Verification
Model Top1/5 Model Top1/5 Model Top1/5
OpenCLIP-L-14 75.2 / 94.3 FaceCLIP-L-14 45.8 / 64.2 FaceCLIP-L-14 75.3 / 94.9
OpenCLIP-bigG-14 80.1 / 96.0 FaceCLIP-bigG-14 47.2 / 69.9 FaceCLIP-bigG-14 76.9 / 95.1
@@ 4.4.1 ID Preservation Comparison
4.5 Ablation Study
In this section, we investigate the impact of the ID alignment and text alignment losses—$\mathcal{L}_{c}(e_{c\to r},e_{r_{\text{cls}}})$ and $\mathcal{L}_{c}(e_{c\to t},e_{t_{\text{cls}}})$, respectively—on the model’s ability to preserve identity, align with text prompts, and maintain image quality. We train three variants of the FaceCLIP encoder using different loss configurations, as indicated in the first column of Table 3. $\mathcal{L}_{1}$ corresponds to the original text-image alignment loss proposed in CLIP [42], which imparts basic text-to-image generation capability to the model. $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ further incorporate alignment with the facial embedding space and text embedding space, respectively. As shown in Table 3, the FaceCLIP encoder trained solely with $\mathcal{L}_{1}$ (CLIP loss) fails to provide identity preservation in the generative model. In contrast, the encoder trained with $\mathcal{L}_{2}$, which combines CLIP loss and ID alignment loss, enables effective identity preservation. Furthermore, as demonstrated in the third row of Table 3, incorporating the text alignment loss $\mathcal{L}_{c}(e_{c\to t},e_{t_{\text{cls}}})$ introduces beneficial regularization, enhancing the model’s text alignment performance.
Pre-training Loss Face Sim. $\uparrow$ CLIP-T $\uparrow$ FID $\downarrow$
$\mathcal{L}_{1}=\mathcal{L}_{c}(e_{c\to t},e_{I})$ 0.077 0.351 86.9
$\mathcal{L}_{2}=\mathcal{L}_{c}(e_{c\to t},e_{I})+\mathcal{L}_{c}(e_{c\to r},e_{r_{\text{cls}}})$ 0.869 0.316 86.9
$\mathcal{L}_{3}=\mathcal{L}_{c}(e_{c\to t},e_{I})+\mathcal{L}_{c}(e_{c\to r},e_{r_{\text{cls}}})+\mathcal{L}_{c}(e_{c\to t},e_{t_{\text{cls}}})$ 0.869 0.332 86.9
Table 3: Impact of different pre-training losses on the final generation quality. The first column lists the loss functions used during the pre-training stage, while the remaining columns report the performance of FaceCLIP-SDXL on various visual metrics.